OpenStudy (anonymous):

Show an equation and a solution for the problem. How much water must be added to 12 L (liters) of a 40% solution of alcohol to obtain a 30% solution? help

OpenStudy (anonymous):

12 L of 40% (by volume) of alcohol solution has 0.4*12=4.8 L of pure alcohol. To make 30% alcohol solution you need to add w L of water. The quantity w can be evaluated with the relation: 4.8 L/(4.8L+w)L=0.3 w=11.2 L

OpenStudy (anonymous):

Show an equation and a solution. How much pure antifreeze must be added to 12 L of a 40% solution to obtain a 60% solution. (Remember that pure antifreeze is 100% = 1.) Question 7 answers what about this

OpenStudy (anonymous):

Ok, you can work this one out yourself. 1) Find how much pure antifreeze you have in the 40% solution. 2) Find how much water needs to be added to make 60% solution.

OpenStudy (anonymous):

than what @cblrtopas

OpenStudy (anonymous):

12*0.40=4.8 L of antifreeze 0.60=(4.8+a)L/12 L --> a=2.4 L of antifreeze must be added.

OpenStudy (anonymous):

Write an equation and a solution for the problem. A health food store sells oatmeal for $3.60 per pound and bran flakes for $4.80 per pound. How many pounds of each should be used to get a mixture of 30 pounds that sells for $4.00 a pound? Question 8 answers i did it already not this one

OpenStudy (anonymous):

The general formula is percent by volume = volume of solute/total volume Just fill in for each quantity placing a variable for the unknown and solve.

OpenStudy (anonymous):

(30 lb)*($4.00/lb)=x*($3.60/lb)+y*($4.80/lb) 120=3.60 x + 4.80 y <--This is equation #1 30 lb = x+y <--This is equation #2 You have two equations and two unknowns; solve the system of equations. 30=x+y --> x=30-y plug this for x in eq.#1 to eliminate it (elimination method). 120=3.60*(30-y) + 4.80 y --> 120 = 108+1.2 y --> y=10 lb Knowing how much y (bran flakes) you need, you can find x (raisins or whatever). 30=x+10 --> x=20 lb